# MCA-Bench: A Multimodal Benchmark for Evaluating CAPTCHA Robustness Against VLM-based Attacks

### v2

@@ 3.3 Data Annotation Strategy
To ensure consistent training and evaluation of multimodal CAPTCHA tasks, we developed a task-driven annotation framework with standardized labeling for four representative task types.
3.3.1 Text-based Tasks: Semantic Consistency Annotation Based on Q&A Structure
We propose a semi-automated, LLM-driven pipeline for generating and filtering math and commonsense questions in text-based CAPTCHA tasks, significantly reducing manual workload while maintaining question quality and diversity. An adaptive prompt template guides the Qwen LLM to produce well-structured and semantically coherent question–answer pairs. Unlike traditional approaches relying on random sampling or manual composition, our method integrates knowledge-constrained generation and task-guided sampling, with type-control parameters to regulate structural complexity, difficulty, and logical patterns.
Following generation, a manual filtering stage applies a multi-dimensional evaluation protocol—assessing grammatical correctness, reasoning consistency, and task ambiguity—to ensure that questions are both discriminative and robust for real-world CAPTCHA evaluation.
3.3.2 Unified Intent Modeling for Coordinate Pointing and Grid Selection
@@
For free-form coordinate pointing tasks, we use a box-to-center projection strategy. Annotators mark the top-left and bottom-right corners of the target’s bounding box using absolute pixel coordinates, with (0, 0) at the top-left of the image. The geometric center of each box serves as the training supervision signal. During inference, a prediction is considered correct if it falls within the corresponding bounding box, following an IoB-Gated Validation mechanism. This approach offers spatial tolerance, improves robustness to outliers, and enhances training convergence.
Formally, the validation criterion between the predicted point $p$ and the bounding box $b=[\,b_{\min},\,b_{\max}\,]$ is defined as:
@@
Here, $\mathbb{I}(\cdot)$ is the indicator function (1 if the condition holds, 0 otherwise), $D=\operatorname{diag}(\,{b}_{\max}-\,{b}_{\min})$ is the diagonal matrix of bounding-box width and height, and $\lVert\cdot\rVert_{\infty}$ is the $\ell_{\infty}$ norm (the maximum absolute component). Thus, the expression tests whether the predicted point lies within the normalized box region centered on the ground-truth box.
We adopt a Discrete Region Index Supervision scheme for grid-based pointing, where each image is divided into a uniform 3×3 grid, yielding nine equal-sized cells indexed from 0 to 8 in row-major order. Annotators label the target region by selecting the corresponding grid index, which serves as a discrete classification target during training. This Semantic-Aligned Index Encoding effectively aligns linguistic cues with local image regions, simplifies supervision under multi-object interference, and enhances model robustness and generalization in the face of linguistic ambiguity and similar visual distractors.
3.3.3 Static Visual Recognition Class Task Labeling Methods
We developed a refined labeling system for static visual recognition tasks to address task diversity and challenges. For character recognition, we use character-level text tagging, where each character in the image is directly labeled and aligned with the content for accurate learning. Labels are automatically generated and validated to prevent errors, with random sampling and manual review to ensure readability and consistency.
For semantic tasks, such as mathematical expressions or symbolic combinations, we apply Visual-Symbolic Abstraction Parsing, parsing the image’s semantic structure into a "final semantic solution". The labels are deterministic and unique, supporting logical reasoning and computation. This system ensures clear, consistent, and task-aligned supervision, providing precise signals for multimodal models and reliable evaluation in complex scenarios.
3.3.4 Interactive Behavior Task Labeling Methods
The annotation system for interaction-based CAPTCHAs focuses on modeling user behavior through detailed trajectory data, target operations, and motion features across common interaction modes such as sliding, rotation, and layer alignment. These tasks combine visual restoration (e.g., aligning puzzle pieces or rotating images) with behavioral trajectory monitoring, allowing the system to distinguish between natural user actions and scripted or emulated inputs. Real user behaviors show non-linearity, irregular speed, and realistic timing, unlike the overly uniform patterns of bots or automation tools.
In slider CAPTCHAs, users drag a puzzle piece into a matching gap; in rotation-based ones, they adjust the image angle via a slider; and in alignment CAPTCHAs, they slide misaligned layers back into position. During these actions, 10 monitoring points capture behavioral data including position, time, and event type. These are analyzed for trajectory naturalness, speed/acceleration profiles, and duration plausibility (data fields detailed in Appendix). For rotation tasks, there is a mapping between the slider’s horizontal displacement ($\Delta x$) and the counter-clockwise rotation angle ($\phi$) of the image. This mapping incorporates affine transformations and periodic functions. Given a maximum slider travel distance of $L=250$ pixels, the rotation angle $\phi$ can be computed as follows:
@@ 3.3.4 Interactive Behavior Task Labeling Methods
[…]
Figure 3: The schematic diagram illustrates the data flow and key module configuration across the four stages of the end-to-end framework: unified interface access, gent fine-tuning loading, collaborative inference execution, structured result feedback.
4 Agent Pipeline and Experiments
@@
Set up We use Qwen2.5-VL-7B-Instruct as the backbone and fine-tune LoRA adapters on 20 CAPTCHA tasks. Images are resized to 224×224, with inputs paired to structured prompts and outputs. Training runs on 4 H20 GPUs with per-GPU batch size 8 and gradient accumulation of 4 (effective batch size 32). We optimize using AdamW with a linearly decaying learning rate starting at $1\times 10^{-4}$. Checkpoints are saved every 100 steps, with early stopping triggered if validation loss fails to improve over 20 evaluation intervals.
MCA-Bench is a high-quality dataset designed to enhance a model’s cracking capabilities We fine-tuned the QWen-2.5VL-7B-instruct model using MCA-Bench, significantly improving its performance. As shown in Figure 4, the fine-tuned model outperforms both the zero-shot baseline and the best closed-source model, ChatGPT-4o, on various CAPTCHA tasks. The dataset boosts the model’s adaptability in visual recognition, logical reasoning, and interaction tasks, while also improving its robustness against complex CAPTCHAs, proving its value as a powerful resource for improving model performance.
Figure 4: Violin plots of Passing-Rate (0–1) for Original Qwen-2.5 VL-7B-Instruct, Original ChatGPT-4-o, and Finetuned Qwen-2.5 VL-7B-Instruct on 20 CAPTCHA classes. Outlines show density; boxes, dots, and whiskers mark IQR, median, and range.
@@
MCA-Bench Dataset Overview MCA-Bench is the first multimodal dataset offering a wide range of CAPTCHA tasks for research. It covers visual recognition, point selection, textual reasoning, and interactive operations, with samples of varying image sizes and features. The training set includes a large number of samples, supporting large-scale training and improving model generalization. The dataset also allows multi-image fine-tuning, optimizing the model’s performance across related tasks. Some tasks use publicly available resources, ensuring diversity and reproducibility in research. Table 1 summarizes key details of the dataset, including categories, image sizes, and training quantities.
Agent System Pipeline The system uses an end-to-end pipeline to unify images, text prompts, and user interactions into a single JSON payload. This payload is directly mapped to a semantic space without intermediate parsing. Based on the task ID in the JSON header, the pipeline activates the appropriate LoRA-Agent, each applying fine-tuned parameters to a shared frozen backbone for efficient multi-task operation with low memory overhead. During inference, visual and textual embeddings are routed to the corresponding agent, which produces outputs such as coordinates, text, mouse movements, or character recognition. These are reassembled into the original JSON format for downstream processing, enabling flexible handling of diverse CAPTCHA tasks, and the overall fine-tuning flow is shown in Figure 3.
Fine-Tuned Model Performance in Visual Tasks For simple visual tasks with minimal background interference, the fine-tuned visual-language model outperforms humans in both speed and accuracy, especially in tasks like target recognition and geometric shape identification. However, as task complexity increases—such as with distorted or rotated images—the model’s performance drops significantly, while humans maintain strong performance in these challenging scenarios. This highlights the limitations of visual-language models in handling complex visual transformations.
@@
Human vs. Model Performance in Reasoning and Interaction Tasks In tasks requiring reasoning or behavioral interaction, the fine-tuned model still lags behind humans. Interactive tasks, such as sliding puzzles and rotating tiles, require precision and coordination, which are challenging for the model. Similarly, in tasks demanding multi-step reasoning or common-sense judgment, the model struggles with complex tasks that involve deeper understanding and contextual reasoning. As shown in Table 2, despite the fine-tuned Qwen-Agent’s success in some standard visual recognition tasks, humans still outperform the model in complex tasks involving reasoning and interaction.
5 Design Principles for Next-Generation Cognitive-Secure CAPTCHAs
Large-scale experiments on MCA-Bench demonstrate that fine-tuned multimodal models achieve over 96% success on visual or shallow-semantic CAPTCHAs but drop below 2.5% for tasks requiring physical interaction or multi-step reasoning. This emphasizes the need for a more robust verification approach, as single-dimension obfuscation is no longer effective against vision–language model-based attacks. To address this, we propose a framework that tightly integrates perception, action, and reasoning through three core principles: Deep Modality Coupling, Behavior-Anchored Validation, and Session-Specific Semantic Personalization.
Deep Modality Coupling involves a continuous challenge pipeline that blends visual cue localization, logical inference, and interactive input. This approach dynamically adjusts task complexity based on user performance, enhancing both security and user experience. Behavior-Anchored Validation analyzes user interaction trajectories, granting verification based on similarity to genuine human data, which prevents scripted attacks without hindering user flow. Session-Specific Semantic Personalization embeds a unique semantic seed into each CAPTCHA, ensuring that solutions rely on session-specific context, preventing automated solvers from precomputing answers. Together, these principles create a cognitively rich verification system that is resilient to advanced attacks.
@@ Conclusion
We present MCA-Bench, the first unified benchmark for evaluating CAPTCHA robustness across twenty real-world tasks spanning visual recognition, text comprehension, logical reasoning, and interactive manipulation. Through extensive experiments with LoRA-based attack agents on state-of-the-art vision–language models, we identify structural vulnerabilities in visual and shallow semantic CAPTCHAs, while tasks involving physical interaction and multi-step reasoning demonstrate greater resilience. Based on these findings, we propose three practical design principles—Deep Modality Coupling, Behavior-Anchored Validation, and Session-Specific Semantic Personalization—to guide the development of more secure human–machine verification systems. MCA-Bench establishes a standardized foundation for adversarial evaluation and future research into adaptive difficulty, human behavior modeling, and usability-aware CAPTCHA design.
@@
A More Details on MCA-Bench A
A.1 Overview of Data Sets A.1
A.2 Distribution Pattern and Training Balance Details for Each Type of Task A.2
A.3 Image Resolution Hierarchy A.3
B Market Research on Mainstream CAPTCHA Datasets B
B.1 Comparative Analysis of User Experience and Security Performance B.1
B.2 Evaluation of Mainstream CAPTCHA Types B.2
B.3 Application Trends and User Acceptance Analysis of Mainstream CAPTCHA Types B.3
B.4 Evaluation and Analysis of Mainstream CAPTCHA Types Selection B.4
B.5 Analysis of CAPTCHA Error-Proneness and User Self-Recovery B.5
B.6 Semiotic Analysis of CAPTCHA Types B.6
C More Extensive Experiments C
C.1 Comparison of Model Stability and Human Performance Variability C.1
C.2 Error-Correction Capability Analysis C.2
C.3 Comparative Analysis of Micro-Decision Path Interpretability C.3
D Limitations and Future Work D
D.1 Limitations D.1
D.2 Future Work Outlook D.2
@@ Appendix A More Details on MCA-Bench
A.1 Overview of Data Sets
To systematically compare the adaptability and data support of different CAPTCHA tasks in model training, this study categorizes them into four major types: static visual recognition, point-based localization, incremental interaction, and textual logical reasoning. We summarize seven key technical indicators for representative datasets, including image resolution, data volume, training set size, original annotation attributes, and whether single-image tasks are supported. Character-based CAPTCHAs typically feature large data volumes and low resolutions, providing abundant and low-cost training samples for deep learning models. In contrast, point-click and interaction-based tasks have relatively smaller datasets but higher annotation complexity, often utilizing high-resolution images that significantly enhance models’ fine-grained representation capabilities in multimodal reasoning tasks such as object localization and action prediction. Interactive CAPTCHAs rely on dynamic generation mechanisms, simulating temporal dependencies and adversarial human-machine interactions in real-world scenarios. For these interactive CAPTCHAs, we have specifically designed a sampling field, as detailed in Table 3.
@@ B.1 Comparative Analysis of User Experience and Security Performance
B.2 Evaluation of Mainstream CAPTCHA Types
@@ B.3 Application Trends and User Acceptance Analysis of Mainstream CAPTCHA Types
B.4 Evaluation and Analysis of Mainstream CAPTCHA Types Selection
To comprehensively evaluate the applicability of different CAPTCHA types in real-world scenarios, this study presents a "Website CAPTCHA Selection Decision Table," constructed based on six core dimensions: security requirements, user experience, development cost, device compatibility, accessibility support, and anti-cracking update frequency. Five mainstream CAPTCHA types are analyzed and quantitatively scored using a five-point scale, where a higher score indicates better performance in the corresponding dimension, as shown in Table 7.
@@ B.4 Evaluation and Analysis of Mainstream CAPTCHA Types Selection
B.5 Analysis of CAPTCHA Error-Proneness and User Self-Recovery
@@
B.6 Semiotic Analysis of CAPTCHA Types
@@
Text-based CAPTCHAs exhibit the highest levels of semantic transparency and signifier-signified correspondence, with a moderate cognitive load. However, their structurally regular nature makes them vulnerable to OCR-based attacks. Image-based CAPTCHAs, benefiting from intuitive visual mappings, demonstrate high schema compatibility across multilingual contexts, achieving a trade-off between usability and security. Behavioral CAPTCHAs (such as sliding puzzles and nine-grid clicks), while lower in semantic transparency and symbolic clarity, require users to quickly learn interaction rules. Nevertheless, their reliance on dynamic trajectories and spatiotemporal uncertainty grants them significant advantages in resisting symbolic deconstruction. Notably, the nine-grid click CAPTCHA—combining visual targeting with spatial interaction—is considered a "2.5D" hybrid form.

### v6

@@ Introduction
We built a standardized, task-driven annotation framework with four representative task types to ensure consistent multimodal CAPTCHA training and evaluation.
Unified Intent Modeling for Coordinate Pointing and Grid Selection
@@
For free-form coordinate pointing tasks, we adopt a box-to-center projection strategy. Annotators label each target by marking the top-left and bottom-right corners of its bounding box using absolute pixel coordinates, with $(0,0)$ at the image’s top-left. The geometric center of the box serves as the training target. During inference, a prediction is considered correct if it falls within the box, following an IoB-Gated Validation rule. This provides spatial tolerance, improves robustness to outliers, and stabilizes training. Formally, the validation criterion for a predicted point $p$ and bounding box $b=[\,b_{\min},b_{\max}\,]$ is defined as:
@@
Here, $\mathbb{I}(\cdot)$ is the indicator function (1 if the condition holds, 0 otherwise), $D=\operatorname{diag}(\,{b}_{\max}-\,{b}_{\min})$ is the diagonal matrix of bounding‐box width and height, and $\lVert\cdot\rVert_{\infty}$ is the $\ell_{\infty}$ norm (the maximum absolute component). Thus, the expression tests whether the predicted point lies within the normalized box region centered on the ground‐truth box.
We propose a discrete region‐index supervision scheme: each image is split into a 3×3 row-major grid, and the target is classified by its cell index, aligning language to local regions and improving robustness under multi-object interference and ambiguous or similar visual cues.
Static Visual Recognition Class Task Labeling Methods
A unified labeling framework adapts to any static-vision task with task-specific supervision: for character recognition, we generate precise, content-aligned tags via automated text generation plus human review to balance speed and accuracy; for higher-level semantics, we use Visual-Symbolic Abstraction Parsing to deterministically translate visual structure into “final semantic solutions,” enabling rigorous logical reasoning and computation—yielding clear, consistent labels that drive precise model training and reliable evaluation even in the most complex cases.
Interactive Behavior Task Labeling Methods
We model interaction-based CAPTCHAs by capturing fine-grained user trajectories, target actions, and motion cues—whether sliding, rotating, or aligning layers. By fusing visual restoration with behavioral analysis, the system flags genuine human input (characterized by nonlinear paths, variable speed, and natural timing) and filters out the uniform, mechanically precise patterns typical of bots.
In slider CAPTCHAs, the user drags a puzzle piece to its gap; in rotation CAPTCHAs, they turn the image with a slider; in alignment CAPTCHAs, they slide layers until they coincide. Ten checkpoints record, and we score each trace for path smoothness, velocity profile, and plausible duration. For rotation, the slider’s horizontal travel $\Delta x\in[0,L]$ ($L=250$ px) maps to the counter-clockwise angle and can be computed as follows:
@@
$\Delta x\in[0,L]$ is the slider’s horizontal displacement in pixels (with L=250L=250\,px), $i^{2}=-1$ the imaginary unit; $\exp\bigl(i\cdot 2\pi\cdot\frac{\Delta x}{L})$ maps this displacement onto the complex unit circle, $\arg(\cdot)$ extracts its phase (in radians), and the factor $\frac{360^{\circ}}{2\pi}$ converts that phase into the rotation angle $\phi(\Delta x)$.
Agent Pipeline and Experiments
@@
We fine‑tune LoRA adapters on Qwen2.5‑VL‑7B‑Instruct across 20 CAPTCHA tasks (hu2022lora). Inputs are 224×224 images paired with structured prompts. Training runs on 4 H20 GPUs (batch size 8, gradient accumulation 4 → effective 32) using AdamW with a linear LR decay from from $1\times 10^{-4}$. We save checkpoints every 100 steps and stop early if validation loss doesn’t improve over 20 evaluations.
Dataset Qwen2.5-VL-7B Human
Point-and-click localization
3×3 grid selection 0.960 0.880
Inverted-letter selection 0.520 0.940
Geometric-shape recognition 0.960 0.980
Brightness discrimination 0.665 0.780
Hollow-pattern recognition 0.995 0.985
Sequential-letter ordering 0.925 0.980
Full-image grid selection 0.350 0.740
Color discrimination 0.990 0.885
Vowel selection 0.975 0.805
Arithmetic selection 0.025 0.780
Rotated-letter selection 0.335 0.745
3×3 jigsaw-swap selection 0.015 0.805
Static visual recognition
Classic character CAPTCHA 0.320 0.920
Distorted-word CAPTCHA 0.985 0.840
Arithmetic-character CAPTCHA 0.695 0.985
Interactive manipulation
Sliding-block 0.360 0.740
Rotation-block 0.280 0.760
Alignment sliders 0.550 0.720
Textual logic Q&A
Text-based arithmetic 0.985 0.970
Commonsense reasoning 0.970 0.860
Table 2: Comparison of Pass Rates for CAPTCHA Types: Qwen2.5-VL-7B vs. Human Performance.
MCA-Bench: Cracking Capability Dataset
Fine-tuning QWen-2.5VL-7B-instruct on MCA-Bench yields substantial gains. As shown in Fig 4, the adapted model surpasses its zero-shot baseline and closed-source peers across CAPTCHA tasks, enhancing visual recognition, logical reasoning, interaction, and robustness to complex challenges.
MCA-Bench Dataset Overview
@@
MCA-Bench is the first multimodal CAPTCHA dataset covering visual recognition, point selection, textual reasoning, and interactive operations. It features varied image sizes, publicly sourced tasks for diversity and reproducibility, and a large training set that supports multi-image fine-tuning to improve generalization.
Agent System Pipeline
The pipeline packs every image, prompt, and user action into one JSON record. A task-ID in the header activates the matching LoRA agent (hu2022lora), which adds its lightweight adapter to a shared frozen backbone. Visual- and text-embeddings flow through that agent to generate outputs—coordinates, strings, mouse traces, or character codes. These results are written back into the original JSON format for downstream use. This design lets the same backbone handle all CAPTCHA tasks with minimal memory and no intermediate parsers (Fig. 3).
Fine-Tuned Model Performance in Visual Tasks
On simple visual tasks with minimal noise, the fine-tuned VLM surpasses human speed and accuracy. However, its performance sharply declines under complex transformations such as distortion or rotation, highlighting VLMs’ limitations compared to robust human perception.
Zero-Shot vs. Fine-Tuned Performance Gaps
Despite differing architectures and training regimes, zero-shot pass rates vary widely across Qwen2.5-VL, ChatGPT-4o, Seed1.5-VL, and Gemini2.5-Pro (Table 1), yet none match the fluid adaptability of human solvers. Fine-tuning on MCA-Bench consistently improves performance—especially for Qwen2.5-VL—but even the best-tuned agents remain below human robustness on multi-step reasoning or precise interaction tasks. This highlights that, while instruction tuning and backbone advances bring notable gains, human proficiency in nuanced, context-rich challenges is still unmatched.
Human vs. AI in Reasoning and Interaction
@@
In tasks requiring reasoning or behavioral interaction—such as sliding puzzles, rotating tiles, multi-step reasoning, and common-sense judgment—the fine-tuned model still lags behind human performance. As shown in Table 2, despite improvements in standard visual recognition, the fine-tuned Qwen-Agent struggles with tasks demanding deeper understanding, contextual reasoning, and precise coordination.
@@ Conclusion
MCA-Bench is the first unified benchmark showing that although LoRA-based attack agents achieve over 96% accuracy on visual and shallow-semantic CAPTCHAs, their performance drops below 2.5% on tasks involving physical interaction or multi-step reasoning—highlighting the limits of single-dimensional obfuscation. To address this gap, we introduce a tightly integrated human–machine verification paradigm built on three principles: (1) deep modality coupling — jointly fusing visual cue localization, logical inference, and interactive input in an adaptively evolving pipeline; (2) behavior-anchored validation — identifying humans through interaction trajectories such as timing, continuity, and subtle irregularities; and (3) session-specific semantic personalization — injecting unique semantic context into each challenge to prevent reuse or precomputation. Together, these principles create a cognitively rich, adaptive verification framework that substantially strengthens robustness against advanced AI-driven attacks.
MCA-Bench: A Multimodal Benchmark for Evaluating CAPTCHA Robustness Against VLM-based Attacks
@@
A  More Details on MCA-Bench
A.1 Overview of Data Sets
A.3 Image Resolution Hierarchy
B  Market Research on Mainstream CAPTCHA Datasets
B.2 Evaluation of Mainstream CAPTCHA Types
B.4 Evaluation and Analysis of Mainstream CAPTCHA Types Selection
B.5 Analysis of CAPTCHA Error-Proneness and User Self-Recovery
B.6 Semiotic Analysis of CAPTCHA Types
C  More Extensive Experiments
C.1 Comparison of Model Stability and Human Performance Variability
C.2 Error-Correction Capability Analysis
C.3 Comparative Analysis of Micro-Decision Path Interpretability
C.4 Performance Comparison Between MCABench and Traditional OCR
C.5 Quantitative Difficulty Taxonomy Across CAPTCHA Modalities
C.6 Human-Centered Cross-Device Interaction Study for Behavioral Ground-Truth Acquisition
C.7 LoRA Adaptation Parameter Configuration for Efficient Model Fine-Tuning
D  Limitations and Future Work
D.1 Limitations
D.2 Future Work Outlook
@@ Appendix A More Details on MCA-Bench
Overview of Data Sets
To systematically compare the adaptability and data support of different CAPTCHA tasks in model training, this study categorizes them into four major types: static visual recognition, point-based localization, interactive manipulation, and textual logical reasoning. We summarize seven key technical indicators for representative datasets, including image resolution, data volume, training set size, original annotation attributes, and whether single-image tasks are supported, with a comprehensive statistical overview of these datasets provided in Table 3. As shown in the table, character-based CAPTCHAs under static visual recognition typically feature large data volumes (consistently 5000 images) and low resolutions (e.g., 160×60), offering abundant and low-cost training samples for basic visual recognition models. In contrast, point-click localization tasks exhibit greater diversity in scale (4999–6505 images) and complexity, with many including ”Original image” annotations and high-resolution examples (e.g., 672×672 for 3×3 grid selection) that enhance fine-grained representation capabilities.
Interactive manipulation tasks (e.g. sliding block, rotation block) in Table 3 share high resolutions (600×360) and ”Original image” annotations, relying on dynamic generation mechanisms to simulate temporal dependencies and adversarial human-machine interactions, for which we have specifically designed a sampling field, as detailed in Table 4. Textual logic Q&A tasks, meanwhile, involve no image data and have smaller sample sizes (1000 each), focusing on language-based reasoning. Together, these dataset characteristics align with diverse training objectives, from foundational recognition to complex interactive reasoning.
Distribution Pattern and Training Balance Details for Each Type of Task
Type Dataset Image size #Images #Train Original image Single image
[…]
Point-and-click localization $\times3$ grid selection $\times672$ 5500 5300
Inverted-letter selection $\times224$ 5000 4800
Geometric-shape recognition $\times224$ 5898 5698
Brightness discrimination $\times224$ 5377 5177
Hollow-pattern recognition $\times224$ 5000 4800
[…]
Full-image grid selection $\times512$ 5015 4815
Color discrimination $\times224$ 5000 4800
Vowel selection $\times224$ 5322 5122
Arithmetic-selection $\times240$ 5065 4865
Rotated-letter selection $\times224$ 5015 4815
$\times3$ jigsaw-swap selection $\times224$ 6505 6305
[…]
Textual logic Q&A Text-based arithmetic 1000 800
Commonsense reasoning 1000 800
Image Resolution Hierarchy
@@ Appendix B Market Research on Mainstream CAPTCHA Datasets
Comparative Analysis of User Experience and Security Performance
@@
Application Trends and User Acceptance Analysis of Mainstream CAPTCHA Types
@@
Evaluation and Analysis of Mainstream CAPTCHA Types Selection
To comprehensively evaluate the applicability of different CAPTCHA types in real-world scenarios, this study presents a ”Website CAPTCHA Selection Decision Table,” constructed based on six core dimensions: security requirements, user experience, development cost, device compatibility, accessibility support, and anti-cracking update frequency. Five mainstream CAPTCHA types are analyzed and quantitatively scored using a five-point scale, where a higher score indicates better performance in the corresponding dimension, as shown in Table 8.
@@
Analysis of CAPTCHA Error-Proneness and User Self-Recovery
@@
Semiotic Analysis of CAPTCHA Types
@@
Text-based CAPTCHAs exhibit the highest levels of semantic transparency and signifier-signified correspondence, with a moderate cognitive load. However, their structurally regular nature makes them vulnerable to OCR-based attacks. Image-based CAPTCHAs, benefiting from intuitive visual mappings, demonstrate high schema compatibility across multilingual contexts, achieving a trade-off between usability and security. Behavioral CAPTCHAs (such as sliding puzzles and nine-grid clicks), while lower in semantic transparency and symbolic clarity, require users to quickly learn interaction rules. Nevertheless, their reliance on dynamic trajectories and spatiotemporal uncertainty grants them significant advantages in resisting symbolic deconstruction. Notably, the nine-grid click CAPTCHA—combining visual targeting with spatial interaction—is considered a ”2.5D” hybrid form.